OpenStudy (anonymous):

ABCD is a quadrilateral. ADE is a straight line. Work out the value of x. (tbc)

OpenStudy (anonymous):

sum of all four interior angles of a quadrilateral is 360 degree here the three given angles are <A=62 deg, <B=136 deg and <C= 90 degree so the fourth interior angle <ADC of the quadrilateral= 360-[62+136+90]=360-288 = 72 Since ADE is a straight line, <ADC + x = 180 degree -----linear pair => 72 + x = 180 => x = 180 - 72 => x = 108

OpenStudy (anonymous):

You have marked x wrongly as 46 degree x is 108 degree as I hv proved above....